A basketball team sells tickets that cost $10, $20, or, for VIP seats, $30. The team has sold 545 tickets overall. It has sold 163 more $20 tickets than $10 tickets. The total sales are $9770. How many tickets of each kind have been sold?
Answer by richwmiller(17219)

a+b+c=545,
10*a+20*b+30*c=9770,
b=a+163
a-b+0=-163
1,1,1,545
10,20,30,9770
1,-1,0,-163

(165,328,52)